OpenStudy (anonymous):

3x+3x+1+3x+2 Use complete sentences to describe the steps taken to simplify this problem. Make sure you include information about the new common denominator, the final simplified expression, and any restrictions.

OpenStudy (anonymous):

i wrote it wrong hold up please

OpenStudy (anonymous):

\[\frac{ 3 }{ x }+\frac{ 3 }{ x+1 }+\frac{ 3 }{ x+2 }\]

OpenStudy (acxbox22):

\[x \neq 0,-1,-2\]

OpenStudy (anonymous):

First, you should write down all the restrictions (values of x which will make the denominators equal to zero. Second, find a common denominator. (the product of the denominators) Lastly, multiply each term by the common denominator and simplify. (this time, it will not be a rational)

OpenStudy (acxbox22):

i already gave you the restrictions